# DynamicKV: Task-Aware Adaptive KV Cache Compression for Long Context LLMs

### v1

@@ 1 Introduction
We validate our DynamicKV on 16 datasets from LongBench (Bai et al., 2023), demonstrating robust performance across multiple models, including LlaMA-3-8B-instruct (Dubey et al., 2024), qwen-2-7B-instruct (Yang et al., 2024), mistral-7b-chat-v0.2 (Jiang et al., 2023), internlm-2.5-7b-chat-1M (Cai et al., 2024). Our DynamicKV exhibits superior overall effectiveness compared to conventional fixed-pattern methods (Zhang et al., 2024b; Li et al., 2024; Nawrot et al., 2024). Notably, DynamicKV can retain full performance while utilizing only 6.9% of the tokens, and in extreme scenarios, it preserves 90% of the performance with just 1.7% of the tokens. Furthermore, experiments on the Needle in a Haystack benchmark revealed that DynamicKV significantly outperforms state-of-the-art (SOTA) methods.
@@ 4 DynamicKV
Figure 3: Overview of our DynamicKV structure and KV cache compression comparison. Left: input prompt, consisting of three parts: Instruct, Question, and Answer. Middle: DynamicKV structure, which progressively and dynamically updates the KV cache in stages to ensure that the total KV cache size remains within the maximum budget. Right: a comparison between DynamicKV and FullKV, highlighting the efficiency and resource savings achieved by our dynamic updating strategy.
here, pooling helps in understanding the context better and $A_{l,h}$ denotes the attention score for the l-th layer and h-th head. This approach allows us to effectively pool the attention scores, ensuring that key tokens are retained based on their relevance to both the current window and previous context.
@@
here, $bs$ represents the size of retained KV cache across all layers. Next, we design a layer-aware progressive dynamic KV cache compression method. The prefill phase of LLMs involves a hierarchical forward process, where for each layer, we retain a KV cache of length $bs$ when computing $A$. Additionally, every $m$ layers, we perform an update across the current and all previous layers. Specifically, for each layer, we use a top-$k$ strategy to retain the largest $bs$ values from $A_{l}$, where $A_{l}$ represents the attention scores of layer $l$. The formula for this process is as follows:
$A_{l}^{\prime}=\text{TopK}(A_{l},bs)$ (3)
@@
$\text{KV}_{l}^{\prime}=\text{KV}_{l}[A_{l}^{\prime}.\text{indices}]$ (4)
@@
Algorithm 1 DynamicKV in Prefill Phase
2:  Output: Compressed K/V cache ${K}^{c}$, ${V}^{c}$
3:  $bs=(wt-ws)\times r_{max}$
4:  def Update_Buffer_Length($A$, $l$):
5:       ${A}^{gather}$leftarrow cat(([${A}$ for $l$ in (1, $l$)]), 0).view(-1)
7:       Compute the $norm$ of $cnts$, range in (0, 1)
8:       $BL$leftarrow [int(($bs*t$ / max($norm$))) for $t$ in $norm$]
9:       $r$leftarrow sum($BL$) / (($wt-ws$)$*L$)
10:       $BL$leftarrow [int($k/r$) for $k$ in $BL$]
11:       Return $BL$
12:  for $l\leftarrow 1\textbf{ to }L$ do
13:     Compute full KV states ${K}^{s}$, ${V}^{s}$
14:     for $h\leftarrow 1\textbf{ to }H$ do
17:     end for
18:     Append ${A}_{l}$ to ${A}^{b}$ /* current ${A}_{l}$ shape is [$H$, $S$] */
19:     /* calculate current layer buffer KV cache */
23:     /* gradually compress*/
[…]
25:        $Bl$leftarrow Update_Buffer_Length(${A}_{l}$, $l$)
26:        /* update the buffer K/V Cache*/
27:        for $i$leftarrow 1 to $l$ do
30:        end for
31:     end if
32:  end for
33:  Update the K/V Cache ${K}^{c},{V}^{c}$ from ${K}^{b},{V}^{b}$
@@ 5.2 Comparative experiments on LongBench
Table 1: Performance comparison on the LongBench dataset for full KV cache, previous methods (StreamingLLM, H2O, SnapKV, PyramidKV), and our DynamicKV method, with KV cache sizes of 128 and 512, using models including LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.2, QWen2-7B-Instruct, and InternLM. Bold indicates the best performance.
@@ A.1 Model Details
Dataset Source Avg len Metric Language #data

### v2

@@ 1 Introduction
We validate our DynamicKV on 16 datasets from LongBench (Bai et al., 2023), demonstrating robust performance across multiple models, including LlaMA-3-8B-Instruct (Dubey et al., 2024), Qwen-2-7B-Instruct (Yang et al., 2024), Mistral-7b-Instruct-v0.2 (Jiang et al., 2023), InternLM-2.5-7b-Chat-1M (Cai et al., 2024). Our DynamicKV exhibits superior overall effectiveness compared to conventional fixed-pattern methods (Zhang et al., 2024b; Li et al., 2024; Nawrot et al., 2024). Notably, DynamicKV can retain full performance while utilizing only 6.9% of the tokens, and in extreme scenarios, it preserves 90% of the performance with just 1.7% of the tokens. Furthermore, experiments on the Needle in a Haystack benchmark revealed that DynamicKV significantly outperforms state-of-the-art (SOTA) methods.
@@ 4 DynamicKV
where pooling helps better understand the context and $A_{l,h}$ denotes the attention score for the l-th layer and the h-th head. This approach allows us to effectively pool the attention scores, ensuring that key tokens are retained based on their relevance to both the current window and previous context.
Figure 3: Overview of our DynamicKV structure and KV cache compression comparison. Left: Layer-wise KV cache retention mechanism in transformer architectures. Right: Our proposed DynamicKV framework employs stage-wise dynamic updating to maintain KV cache within predefined memory budgets, with task-specific visualization showing KV cache preservation patterns across layers.
@@
where $bs$ represents the size of retained KV cache across all layers. Next, we design a layer-aware progressive dynamic KV cache compression method. The prefill phase of LLMs involves a hierarchical forward process, where for each layer, we retain a KV cache of length $bs$ when computing $A$. Additionally, every $m$ layers, we perform an update across the current and all previous layers. Specifically, for each layer, we use a top-$k$ strategy to retain the largest $bs$ values from $A_{l}$, where $A_{l}$ represents the attention scores of layer $l$. The formula for this process is as follows:
$A_{l}^{\prime}=\text{TopK}(A_{l},bs).$ (3)
@@
$\text{KV}_{l}^{\prime}=\text{KV}_{l}[A_{l}^{\prime}.\text{indices}],$ (4)
@@ 5.2 Comparative experiments on LongBench
Table 1: Performance comparison on the LongBench dataset for full KV cache, previous methods (StreamingLLM, H2O, SnapKV, PyramidKV), and our DynamicKV method, with KV cache sizes of 128 and 512, using models including LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.2, QWen2-7B-Instruct, and InternLM-2.5-Chat-1M. Bold indicates the best performance.
@@ 5.4 Ablation Study
☞ More analyses in Appendix
In addition to the above discussions, we conduct more related analyses and show them in Appendix: (1) DynamicKV is effectively scalable to other models on the Needle-in-Haystack task (Appendix A.3); (2) Enhancing generation efficiency with 22.5-129% higher token throughput compared to fullKV and superior memory optimization saving up to 11.2% GPU memory while maintaining stable performance (Appendix A.4); (3) Task-Aware adaptability that dynamically adjusts token processing across transformer layers based on input complexity(Appendix A.5). Please refer to the Appendix for more details.
@@ 6 Conclusion
Limitations
Our work has several potential limitations. First, given the limited computational budget, we only validate our DynamicKV on models Scaling up to super-large model sizes (e.g., 70B), and applying DynamicKV to more cutting-edge model architectures will be more convincing model architectures. Second, although we have conducted experiments on multiple tasks including single- and multi-document QA, summarization, synthetic tasks, and code completion, the generalization ability of DynamicKV to other tasks or datasets has not been fully explored. Future work will focus on expanding the application scope of DynamicKV to more diverse tasks and datasets.
@@ A.1 Model Details
Dataset Source Avg length Metric Language #data
@@ A.2 Dataset Details
Model StreamingLLM H2O SnapKV PyramidKV DynamicKV
LlaMA-3-8B-Instruct 0.29 0.46 0.80 0.89 0.9
Qwen-2-7B-Instruct 0.22 0.41 0.84 0.86 0.87
Table 5: Configuration of Models.
A.3 Need in a HayStack
The experimental results are presented in Table 5, which illustrates the performance of our DynamicKV approach across various models, including LlaMA and Qwen, in Needle in a Haystack task.
Input Len Output Len Method TTFT (s) TPOT (tok/s) Latency (s) Memory (MB)
8k 2k FullKV 0.66 27.63 74.79 20055
8k 2k Dynamickv 0.70 33.85 61.21 19417
16k 4k FullKV 1.45 19.55 209.56 23859
16k 4k Dynamickv 1.49 33.02 125.52 22051
32k 8k FullKV 3.52 11.65 706.56 31213
32k 8k Dynamickv 3.58 26.69 310.56 27713
Table 6: Efficiency Comparison Between FullKV and DynamicKV
A.4 Efficiency Experiments
We evaluate the efficiency of DynamicKV against the standard method (FullKV) under varying input/output lengths. All experiments are conducted with a fixed context window ($m=128$), measuring Time-to-First-Token (TTFT), Time-Per-Output-Token (TPOT), end-to-end latency, and GPU memory usage. The results are summarized in Table 6.
Key observations include:
  • Short Sequences (8k/2k): DynamicKV improves TPOT by 22.5% (27.63→33.85 tok/s) while slightly increasing TTFT by 6% (0.66s→0.70s), achieving 18.2% lower total latency (74.79s→61.21s) with 638MB memory reduction.
  • Long Sequences (32k/8k): The advantages amplify significantly, with DynamicKV delivering 129% higher TPOT (11.65→26.69 tok/s), 56% lower latency (706.56s→310.56s), and 11.2% memory savings (31213MB→27713MB).
  • Scalability: FullKV shows superlinear TPOT degradation (11.65 tok/s at 32k inputs), while DynamicKV maintains stable throughput through on-demand computation, demonstrating better adaptability to long-context generation.
The experiments demonstrate that dynamic KV caching trades marginal initial latency for substantially better sustained generation speed and memory efficiency, particularly beneficial for long-text generation tasks (>2k output tokens).
Dataset Layer-wise ( 1 - 32 ) Tokens
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32
NarrativeQA 93 75 111 51 71 75 91 108 97 154 138 143 196 126 156 201
173 133 202 200 158 128 115 86 123 109 117 76 122 120 142 206
Qasper 157 66 92 43 59 70 77 111 93 137 132 160 187 122 147 186
184 132 183 181 160 135 122 77 119 117 128 69 127 113 179 231
MultifieldQA-en 245 79 109 50 76 75 92 106 90 169 138 164 205 115 155 196
174 116 167 169 145 116 103 74 97 90 108 67 116 116 154 220
HotpotQA 145 75 115 55 73 73 92 102 88 146 128 140 190 106 140 208
172 129 194 185 146 123 109 80 105 99 125 75 127 122 173 256
2WikiMQA 198 74 110 53 69 74 91 105 89 159 133 153 196 109 143 215
178 132 181 179 139 112 96 76 102 94 120 70 118 117 161 250
MuSiQue 108 71 108 54 72 71 89 98 85 143 124 135 189 104 138 216
178 128 206 198 148 130 113 84 112 100 126 76 134 128 173 257
GovReport 332 98 113 48 73 78 88 136 125 150 176 145 142 113 136 164
212 129 148 161 145 115 79 62 103 78 94 57 99 121 138 238
QMSum 156 85 119 51 65 63 78 116 81 146 126 136 185 105 142 218
187 121 204 201 165 134 121 85 122 105 122 70 115 121 144 207
MultiNews 459 94 115 50 67 71 88 101 99 120 126 147 168 130 156 190
202 125 147 160 134 107 78 60 101 84 95 55 102 117 146 202
TREC 214 70 112 47 69 97 93 129 98 156 142 215 220 141 213 115
143 110 166 103 105 91 90 67 84 89 139 78 129 96 192 283
TriviaQA 105 62 97 52 73 89 86 106 102 178 139 178 198 140 189 199
168 118 159 139 144 103 90 66 109 102 131 71 127 124 196 256
SAMSum 126 64 103 48 71 96 93 152 118 157 130 174 173 150 189 179
168 132 177 135 135 108 98 84 101 104 120 87 114 114 168 228
PassageCount 91 72 106 48 54 58 80 111 96 119 153 143 166 147 157 225
214 147 191 203 178 106 96 72 116 116 118 70 122 130 148 243
PassageRetrieval-en 152 92 129 70 72 75 92 121 108 128 162 149 142 139 145 175
179 145 180 172 134 113 95 87 112 109 124 91 108 118 140 238
LCC 186 59 104 44 64 80 98 128 112 159 169 161 177 173 162 163
175 133 170 147 113 102 83 69 94 80 114 84 138 130 178 247
RepoBench-P 106 63 106 46 66 89 96 132 119 168 177 169 179 183 176 189
197 138 177 148 113 102 80 71 94 80 106 83 132 133 157 221
Table 7: Per-layer token processing statistics across 32 transformer layers by DynamicKV in Mistral-7B-Instruct-v0.2 model on LongBench datasets.
A.5 Task-Aware Sensitivity Study
The presented Table 7 offers a comprehensive insight into the token processing statistics per layer for 32 transformer layers using the DynamicKV method in Mistral-7B-Instruct, as applied to various datasets from LongBench. This detailed analysis reveals significant variability in how tokens are processed at different layers depending on the specific task.
In particular, our approach demonstrates that certain tasks, such as those represented by the MultiNews and GovReport datasets, exhibit exceptionally high token processing counts even in early layers, indicating a potentially greater complexity or information density in these tasks. Conversely, datasets such as TriviaQA and SAMSum show relatively lower initial processing demands, but with significant spikes in later layers, suggesting evolving complexity throughout the document or dialogue.
Moreover, the data highlights that the DynamicKV mechanism adeptly adjusts to the unique demands of each dataset, optimizing performance and resource allocation dynamically across all layers. This adaptability is crucial for handling the diverse range of tasks and input lengths encountered in real-world applications, ensuring efficient processing without compromising on understanding or accuracy.
Algorithm 1 DynamicKV in Prefill Phase
2:  Output: Compressed K/V cache ${K}^{c}$, ${V}^{c}$
3:  $bs=(wt-ws)\times r_{max}$
4:  def Update_Buffer_Length($A$, $l$):
5:       ${A}^{gather}$leftarrow cat(([${A}$ for $l$ in (1, $l$)]), 0).view(-1)
7:       Compute the $norm$ of $cnts$, range in (0, 1)
8:       $BL$leftarrow [int(($bs*t$ / max($norm$))) for $t$ in $norm$]
9:       $r$leftarrow sum($BL$) / (($wt-ws$)$*L$)
10:       $BL$leftarrow [int($k/r$) for $k$ in $BL$]
11:       Return $BL$
12:  for $l\leftarrow 1\textbf{ to }L$ do
13:     Compute full KV states ${K}^{s}$, ${V}^{s}$
14:     for $h\leftarrow 1\textbf{ to }H$ do
17:     end for
18:     Append ${A}_{l}$ to ${A}^{b}$ /* current ${A}_{l}$ shape is [$H$, $S$] */
19:     /* calculate current layer buffer KV cache */
23:     /* gradually compress*/
[…]
25:        $Bl$leftarrow Update_Buffer_Length(${A}_{l}$, $l$)
26:        /* update the buffer K/V Cache*/
27:        for $i$leftarrow 1 to $l$ do
30:        end for
31:     end if
32:  end for
33:  Update the K/V Cache ${K}^{c},{V}^{c}$ from ${K}^{b},{V}^{b}$